Exchange Corp. is a company that acts as a facilitator in tax-favored real estate swaps. Such swaps, know as 1031 exchanges, permit participants to avoid some or all of the capital gains taxes that would otherwise be due. The bookkeeper for the company has been asked to prepare a report for the company to help its owner/manager analyze performance. The first such report appears below:

Exchange Corp.
Analysis of Revenues and Costs
For the Month Ended May 31
Actual Unit Revenues
and Costs
Planning Budget
Unit Revenues
and Costs
Variances
Exchanges completed 50 40
Revenue $ 385 $ 395 $ 10 U
Expenses:
Legal and search fees 184 165 19 U
Office expenses 112 135 23 F
Equipment depreciation 8 10 2 F
Rent 36 45 9 F
Insurance 4 5 1 F
Total expense 344 360 16 F
Net operating income $ 41 $ 35 $ 6 F

Note that the revenues and costs in the above report are unit revenues and costs. For example, the average office expense is $135 per exchange completed on the planning budget; whereas, the average actual office expense is $112 per exchange completed.

Legal and search fees is a variable cost; office expenses is a mixed cost; and equipment depreciation, rent, and insurance are fixed costs. In the planning budget, the fixed component of office expenses was $5,200.

All of the company’s revenues come from fees collected when an exchange is completed.

Required:

1. Is the report prepared by the bookkeeper useful as a performance report?

2. Complete the performance report that would help the owner/manager assess the performance of the company in May.
Answer #1

1. No

2.

Exchange Corp Flexible Budget Performance Report For the Month Ended May 31 Actual Revenue & Spending Flexible Results Varian

Working

Variable Budgeted office expenses per unit =  average office expense per exchange completed on the planning budget - fixed component of office expenses per exchange completed on the planning budget

Variable Budgeted office expenses per unit = 135 - 5200/40

Variable Budgeted office expenses per unit = $ 5

Planning Budget

Revenue = 395*40 = $ 15800

office expenses = 5*40 + 5200 = $ 5400

Legal and search fees = 165*40 = $ 6600

Equipment depreciation = 10*40 = $ 400

Rent = 45*40 = 1800

Insurance = 5*40 = 200

Flexible Budget

Revenue = 395*50 = $ 19750

office expenses = 5*50 + 5200 = $ 5450

Legal and search fees = 165*50 = $ 8250

Equipment depreciation = 10*40 = $ 400

Rent = 45*40 = 1800

Insurance = 5*40 = 200

Actual

Revenue = 385*50 = $ 19250

office expenses = 112*50 = $ 5600

Legal and search fees = 184*50 = $ 9200

Equipment depreciation = 8*50 = $ 400

Rent = 36*50 = 1800

Insurance = 4*50 = 200